Jefferson Lab HALL A Logbook for December 01 - All EXCEPT Automatic Entries

  • Main INDEX

    Monday

    
    

    Sunday

    
    

    Saturday

    
    

    Friday

  • 79548 12/28/01 14:42 atrig Trigger-Download
  • Thursday

    
    

    Wednesday

    
    

    Tuesday

    
    

    Monday

    
    

    Sunday

    
    

    Saturday

    
    

    Friday

    
    

    Thursday

  • 79345 12/20/01 14:09 x. Jiang Fpp low voltage and threshold
  • Wednesday

  • 79339 12/19/01 17:33 M JOnes fpp threshold on
  • Monday

  • 0 12/17/01 10:07 no_name no_key
  • Saturday

    
    

    Monday

  • 79304 12/10/01 07:40 f.cusanno shift summary
  • 79299 12/10/01 05:25 atrig Trigger-Download
  • 79296 12/10/01 04:39 atrig Trigger-Download
  • 79293 12/10/01 03:53 atrig Trigger-Download
  • 79290 12/10/01 03:04 atrig Trigger-Download
  • 79289 12/10/01 03:01 cusanno magnet settings
  • 79285 12/10/01 02:11 atrig Trigger-Download
  • Sunday

  • 79271 12/09/01 23:37 fissum HRSl beta (small peak)
  • 79265 12/09/01 23:00 atrig Trigger-Download
  • 79264 12/09/01 22:50 fissum RICH tests again
  • 79263 12/09/01 22:50 fissum staging for beam
  • 79226 12/09/01 21:56 fissum reboot failed; no CW because no ion chambers
  • 79210 12/09/01 21:13 atrig Trigger-Download
  • 79209 12/09/01 21:10 atrig Trigger-Download
  • 79208 12/09/01 21:01 fissum fissum
  • 79207 12/09/01 20:58 atrig Trigger-Download
  • 79195 12/09/01 18:18 fissum close up; request 10 uA
  • 79183 12/09/01 16:48 x. jiang HV channel 13.0 reset, left arm.
  • 79181 12/09/01 16:08 sarty Running Shift Summary (8-16) - RICH test prep
  • 79171 12/09/01 13:42 sarty Aerogel2 HV settings
  • 79168 12/09/01 12:50 atrig Trigger-Download
  • 79167 12/09/01 12:40 sarty Angle and Momentum Settings for RICH test
  • 79166 12/09/01 12:10 Jack Right Arm, left door
  • 79163 12/09/01 09:44 atrig Trigger-Download
  • 79161 12/09/01 08:11 Iodice shift summary
  • 79130 12/09/01 00:06 Mattias Shift summary
  • Saturday

  • 79127 12/08/01 23:29 Mattias Magnet screen with correct angle
  • 79121 12/08/01 22:42 Mattias Target thickness calibration
  • 79119 12/08/01 22:08 fissum chain analysis king1-
  • 79110 12/08/01 20:00 Mattias Magnet screen for q
  • 79106 12/08/01 19:34 fissum GI kinematics overlap e t c
  • 79105 12/08/01 19:29 Mattias GI- finished
  • 79090 12/08/01 16:29 fissum king2 - more on the knockout ridge
  • 79086 12/08/01 16:01 sarty Running Shift Summary (8-16)
  • 79071 12/08/01 12:43 sarty G1- ChainAnalysis (so far)
  • 79048 12/08/01 08:03 Iodice shift summary
  • 79014 12/08/01 01:08 Iodice Emiss for kin g1-
  • 79012 12/08/01 00:18 fissum king2-: correlated 2N-knockout ridge!!
  • 79008 12/08/01 00:09 Iodice Hydrogen elast. radiat. tail
  • 79006 12/08/01 00:05 x. jiang shift summary
  • Friday

  • 79005 12/07/01 23:45 fissum decide to move to kingI-
  • 78997 12/07/01 22:28 fissum chain analysis king2+ (no caps or shield)
  • 78996 12/07/01 22:17 fissum chain analysis gI+
  • 78985 12/07/01 20:25 x. jiang Coda crashed with ROC1 stuck
  • 78982 12/07/01 19:45 Reitz EP results
  • 78970 12/07/01 17:46 fissum gII phase space continued...
  • 78969 12/07/01 17:29 fissum phase space for gII kinematics
  • 78964 12/07/01 16:41 x. jiang moved OTR 1C12 ans 1H04A in.
  • 78958 12/07/01 16:04 LeRose Shift Summary
  • 78955 12/07/01 15:42 LeRose chainalyze_g-
  • 78937 12/07/01 12:59 Reitz Bull Eye Scan
  • 78923 12/07/01 12:11 Reitz BCM Calibration
  • 78904 12/07/01 07:55 bogdanw shift summary
  • 78862 12/07/01 00:05 sarty Running shift summary (16:00-24:00)
  • Thursday

  • 78861 12/06/01 23:59 fissum first g2- results (plots and explanation)
  • 78860 12/06/01 23:46 fissum summary of "capped" spectrometer runs
  • 78858 12/06/01 23:29 fissum optimized prescale factors, current for 20% CPU deadtime
  • 78853 12/06/01 22:47 fissum coincidence peak for g2-
  • 78846 12/06/01 21:39 fissum chain analysis king2+
  • 78843 12/06/01 21:23 sarty Magnet Screen for kinG2- (plus 2hr beam current history)
  • 78840 12/06/01 20:46 fissum DAQ hangup
  • 78838 12/06/01 19:57 fissum king2+ completed
  • 78833 12/06/01 18:50 fissum g2+ analysis routine
  • 78830 12/06/01 18:36 fissum OTRs reinserted
  • 78828 12/06/01 17:46 Wendy Q vec measurement
  • 78825 12/06/01 16:53 fissum guards left in place - Emiss problem solved
  • 78822 12/06/01 16:00 Eric Voutier Shift Summary
  • 78818 12/06/01 11:24 Wendy G plus and G1 plus setting
  • 78815 12/06/01 07:57 bogdanw Shift's Summary
  • 78805 12/06/01 00:36 Iodice shift summary
  • 78802 12/06/01 00:19 Iodice Gplus vs G1plus vs G2plus
  • Wednesday

  • 78794 12/05/01 23:04 Iodice G+ re-analysis
  • 78789 12/05/01 20:58 iodice hydrogen elastic peak
  • 78788 12/05/01 20:35 Wendy G plus and G1 plus settings
  • 78783 12/05/01 18:28 iodice emiss peaks.
  • 78782 12/05/01 18:27 Wendy Omega vs. Q
  • 78779 12/05/01 18:12 iodice emiss funny peak
  • 78778 12/05/01 18:01 Iodice omega_vs_q band
  • 78773 12/05/01 16:07 Mattias Shift summary
  • 78771 12/05/01 15:35 Wendy Coincidence timing for G1+
  • 78768 12/05/01 15:04 Mattias Magnet screen for G_I_+
  • 78755 12/05/01 08:01 Eric Voutier Shift Summary
  • 78749 12/05/01 06:20 R. Michaels reboot is an alias
  • Tuesday

  • 78734 12/04/01 23:32 Hyde-Wright daq
  • 78728 12/04/01 18:25 Hyde-Wright OLO2 Calibration
  • 78724 12/04/01 13:57 Wendy K plus kinematics
  • 78723 12/04/01 12:11 Saha Hall OTR update from J.C. Denard
  • 78721 12/04/01 07:58 X. Zheng shift summary
  • 78708 12/04/01 00:30 V. Sulkosky Correction on beginning of Runs
  • 78706 12/04/01 00:03 P. Moussiegt shift summary
  • Monday

  • 78703 12/03/01 23:43 fissum kplus analysis routine
  • 78700 12/03/01 22:13 fissum start kplus at 50 uA
  • 78698 12/03/01 20:11 fissum kplus coincidence peak
  • 78697 12/03/01 19:55 fissum kplus timing peak
  • 78696 12/03/01 19:39 fissum kplus detector replay OK
  • 78695 12/03/01 19:37 fissum kin k+ detectors
  • 78692 12/03/01 19:17 fissum chain analysis of jplus data
  • 78691 12/03/01 19:13 fissum recompiled recon
  • 78689 12/03/01 17:27 fissum adjusted prescales for K+
  • 78687 12/03/01 17:14 fissum moved to kin K+
  • 78685 12/03/01 16:19 Wendy Shift Summary
  • 78677 12/03/01 12:09 Wendy J plus above 50 MeV
  • 78676 12/03/01 11:40 Wendy J plus
  • 78673 12/03/01 09:36 Wendy Angle Pin
  • 78671 12/03/01 08:06 X. Zheng shift summary
  • 78667 12/03/01 07:03 R. Michaels adaql1 reboot
  • 78663 12/03/01 05:58 X. Zheng adaql1 not working
  • 78656 12/03/01 00:12 Mattias Shift summary
  • Sunday

  • 78654 12/02/01 23:02 Mattias Analysis from mss cache
  • 78651 12/02/01 22:23 Mattias TC-corr for J+
  • 78650 12/02/01 22:05 fissum detector analysis kin J+
  • 78646 12/02/01 17:34 fissum i- summary
  • 78642 12/02/01 16:23 weinstein shift summary
  • 78639 12/02/01 15:45 Reitz Screen for kin j+
  • 78622 12/02/01 11:31 weinstein waterfall target reboot
  • 78611 12/02/01 09:08 weinstein beam spot size - vertical
  • 78608 12/02/01 08:02 R. Michaels owl shift
  • 78591 12/02/01 05:12 R. Michaels synch checks - ok
  • 78578 12/02/01 01:55 R. Michaels OTR says Y is stable
  • 78574 12/02/01 01:38 R. Michaels fast beam motion in vertical ??
  • 78572 12/02/01 01:00 R. Michaels end-run halog scalers fixed
  • Saturday

  • 78566 12/01/01 23:59 Mattias Shift summary
  • 78557 12/01/01 22:29 Mattias I- chain results
  • 78554 12/01/01 22:16 Mattias RICH transparancy test
  • 78546 12/01/01 20:48 Mattias I- Emiss
  • 78541 12/01/01 19:34 Mattias I- so far
  • 78540 12/01/01 19:31 Mattias CODA crash
  • 78529 12/01/01 17:17 Mattias Left Q2 tripped
  • 78524 12/01/01 16:21 Wendy Shift summary
  • 78497 12/01/01 10:35 Wendy Current Spikes
  • 78483 12/01/01 08:07 A. Camsonne Owl shift summary
  • 78475 12/01/01 06:38 A. Camsonne Waterfall target IOC reboot
  • 78472 12/01/01 06:16 A. Camsonne Target controls MEDM EPICS
  • 78452 12/01/01 01:59 A. Camsonne Beam off : QE measurement
  • 78446 12/01/01 01:12 A. Camsonne Run 2716 CODA crashed
  • 78442 12/01/01 00:04 Mattias Shift summary